# Annabelle Michael Carrell    Albert Gong    Abhishek Shetty    Raaz Dwivedi    Lester Mackey

### v5

@@ 4.2 Thinning attention in practice
To gauge the practical effectiveness of Alg. 1, we recreate the benchmark Tokens-To-Token Vision Transformer (T2T-ViT) experiment of Zandieh et al. (2023). In this experiment, attention approximations are scored on their ImageNet classification accuracy and computational expense when used as drop-in replacements for the two most expensive attention layers in a pretrained T2T-ViT neural network (Yuan et al., 2021). Using the exact implementations and settings provided by Zandieh et al. (2023), we benchmark our PyTorch implementation of Thinformer against exact attention and four leading attention approximations: Performer (Choromanski et al., 2021), Reformer (Kitaev et al., 2020), ScatterBrain (Chen et al., 2021), and KDEformer. In Tab. 3, we find that Thinformer provides the highest Top-$1$ accuracy on the ImageNet 2012 validation set (Russakovsky et al., 2015), while running faster than all of the alternatives. The final attention call of Thinformer can also be combined with optimized attention implementations like FlashAttention (Dao et al., 2022; Dao, 2024) to further reduce the time and memory footprint. We provide PyTorch code replicating this experiment at https://github.com/microsoft/thinformer and supplementary experiment details in Sec. L.1.
@@ 5 Faster SGD Training
To train a machine learning model parameterized by a vector $\boldsymbol{w}\in\mathbb{R}^{d}$, a standard approach is to minimize the empirical risk $f(\boldsymbol{w})\triangleq\frac{1}{n}\sum_{i=1}^{n}f_{i}(\boldsymbol{w})$ using stochastic gradient descent (SGD) updates,
@@
for each epoch $k\in[K]$ and datapoint $i\in[n]$. Here, $\alpha>0$ is a step size, each $f_{i}$ is a datapoint-specific loss function, and $\pi_{k}$ is a permutation of $[n]$ representing the order in which datapoints are processed in the $k$-th epoch.
@@
Figure 1: Train and test convergence trajectories for mortgage classification with reordered SGD variants. We display mean values ±$\pm 1$ standard deviation across $5$ random seeds. See Sec. 5.2 for more details.
@@ 5.1 Bridging the dimension gap
$\|{\nabla f_{i}(\boldsymbol{w})-\nabla f(\boldsymbol{w})}\|_{2}^{2}\leq\sigma^{2},$ (30)
$\|{\nabla f_{i}(\boldsymbol{w})-\nabla f_{i}(\boldsymbol{v})}\|_{2}\leq L\|{\boldsymbol{w}-\boldsymbol{v}}\|_{2},\quad\text{and}\quad$ (31)
$f(\boldsymbol{w})-f^{\star}\leq\frac{1}{2\mu}\|{\nabla f(\boldsymbol{w})}\|_{2}^{2}\quad\text{for}\quad f^{\star}\triangleq\inf_{\boldsymbol{v}\in\mathbb{R}^{d}}f(\boldsymbol{v}).$ (32)
Then SGD 29 with LKH$(\frac{1}{2K})$ reordering (Alg. 2) and step size $\alpha$ given in App. G satisfies, with probability at least $\frac{1}{2}$,
$\textstyle f(\boldsymbol{w}_{K})-f^{\star}\leq\widetilde{O}(\frac{r}{n^{2}K^{2}})\quad\text{for}\quad\qquad\qquad$ (33)
$r\triangleq\max_{k\in[K]}\mathrm{rank}_{\epsilon_{k}}(\mathbf{X}^{k}),\quad\mathbf{X}^{k}\triangleq[\boldsymbol{x}_{1}^{k},\dots,\boldsymbol{x}_{n}^{k}]^{\top},\quad$ (34)
$\textstyle\epsilon_{k}\triangleq{\max_{i\in[n]}\|{\boldsymbol{x}_{i}^{k}-\bar{\boldsymbol{x}}^{k}}\|_{2}}/{\sqrt{n}},\text{\ \ and\ \ }\bar{\boldsymbol{x}}^{k}\triangleq\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{x}_{i}^{k}.$ (35)
The proof of Thm. 3 in App. G simply uses Thm. 1 to bound the thinning quality of LKH$(\frac{1}{2K})$ and then adapts the prior SGD analysis of Cooper et al. (2023). Notably, the standard practice of random reshuffling, i.e., SGD with uniform reordering, can only guarantee a significantly slower $\Omega(\frac{1}{nK^{2}})$ rate under these assumptions (Rajput et al., 2020, Thm. 2), while Lu et al. (2022, Thm. 4) implies a similar but dimension-dependent $\widetilde{O}(\frac{d}{n^{2}K^{2}})$ rate for SBW reordering. Thm. 3 shows that this dimension dependence can be avoided whenever the gradient update matrices $\mathbf{X}^{k}$ are low-rank, or, more generally, whenever they are $\epsilon=O(1/\sqrt{n})$-approximable by low-rank matrices.
@@ 5.2 Bridging the theory-practice gap
Two criticisms levied by Lu et al. (2022) against the SBW algorithm were the need to estimate the maximum Euclidean norm of any possible gradient vector in advance and the need to tune its free hyperparameter. LKH$(\frac{1}{2K})$ has neither of these drawbacks as it automatically adapts to the scale of each input and has no hyperparameters to tune. Moreover, with a linear kernel, LKH$(\frac{1}{2K})$ can be run online in $O(nd)$ time. Hence, LKH$(\frac{1}{2K})$ is a promising substitute for the greedy thinning of Lu et al. (2022); Cooper et al. (2023). Indeed, when we recreate the Home Mortgage Disclosure Act logistic regression experiment of Cooper et al. (2023) with a single worker (Fig. 1), we find that LKH-SGD strongly outperforms the standard practice of random reshuffling (RR) and the theoretically justified but overly conservative CD-GraB: SBW variant. In addition, LKH-SGD matches the state-of-the-art test accuracy of CD-GraB: Greedy and lags only slightly in terms of training convergence. See https://github.com/microsoft/khsgd for PyTorch code replicating this experiment and Sec. L.2 for supplementary experiment details.
@@ 6 Cheap Two-Sample Testing
A core task in statistics and machine learning is to determine whether two datasets are drawn from the same underlying distribution. In this two-sample testing problem, we observe independent samples $\mathcal{X}\triangleq(\boldsymbol{x}_{i})_{i=1}^{m}$ and $\mathcal{Y}\triangleq(\boldsymbol{y}_{j})_{j=1}^{n}$ from the unknown distributions $\mathbb{P}$ and $\mathbb{Q}$ respectively, and we seek to accept or reject the null hypothesis that $\mathbb{P}=\mathbb{Q}$. Standard kernel MMD tests tackle this task by computing the empirical MMD
@@ 6.1 Low-rank analysis of Compress Then Test
$O(4^{\mathfrak{g}}(m+n)(s+\log_{4}\mathopen{}\mathclose{{}\left(\frac{m+n}{s}-\mathfrak{g}}\right))),$ (38)
@@
Instantiate the assumptions of Thm. 4 with $\mathbf{k}=\mathbf{k}_{\textup{deep}}$ 43. If the inputs $(\phi(\boldsymbol{x}_{1}),\boldsymbol{x}_{1},\phi(\boldsymbol{y}_{1}),\boldsymbol{y}_{1})$ are sub-Gaussian, that is,
@@
Cors. 3 and 4 follow from explicitly bounding the eigenvalues of the generated deep kernel matrices as in 17 and 20. One could alternatively bound the compression error of KT-Compress$(\delta)$ using the covering number approach of Dwivedi & Mackey (2022, Thm. 2, Prop. 3). In the setting of Cor. 3, the argument of App. J combined with this distinct analysis would yield an alternative error inflation factor $\mathbsf{\tilde{R}}_{\mathbf{k}_{\textup{deep}}}/2^{\mathfrak{g}}$ with worse dimension dependence,
@@ 6.2 Powerful deep kernel testing in near-linear time
Figure 2: Time-power trade-off curves for detecting Higgs bosons with deep kernel MMD tests. We plot mean values ±$\pm 1$ standard error across $1000$ independent trials with level $\alpha=0.05$ and $\mathcal{B}=100$ permutations.
@@
Fig. 2 compares the time-power trade-off curves induced by three fast kernel testing approaches to this problem: Subsampling, a standard wild-bootstrap MMD test (Chwialkowski et al., 2014) that simply evaluates empirical $\operatorname{MMD}_{\mathbf{k}_{\textup{deep}}}$ using $n_{\textup{out}}=m_{\textup{out}}$ uniformly subsampled points; W-Block, a wild-bootstrap test that averages $\frac{n}{B}$ subsampled squared $\operatorname{MMD}_{\mathbf{k}_{\textup{deep}}}$ estimates based on $n_{\textup{out}}=m_{\textup{out}}=B$ points (Zaremba et al., 2013); and CTT with $s=32$ bins and varying $\mathfrak{g}$. We find that the CTT curve dominates that of the alternative methods and matches the power of an exact MMD test (Subsampling with $n_{\textup{out}}=n$) in a fraction of the time. See https://github.com/microsoft/deepctt for PyTorch code replicating this experiment and Sec. L.3 for supplementary experiment details.
@@ Appendix A Appendix Notation and Definitions
If, in addition, the event has probability $1$, we say that $\boldsymbol{w}$ is $(\mathbf{K},\nu)$-sub-Gaussian.
Notably, a thinning algorithm is $(\mathbf{K},\nu,\delta)$-sub-Gaussian if and only if its associated vector $\boldsymbol{p}_{\textup{in}}-\boldsymbol{p}_{\textup{out}}$ is $(\mathbf{K},\nu)$-sub-Gaussian on an event $\mathcal{E}$ of probability at least $21-\delta/$ 2.
@@
If, in addition, the event has probability $1$, we say that $\phi$ is $(\mathbf{k},\nu)$-sub-Gaussian.
@@
Invoking the representations 52 and the functional sub-Gaussianity condition 50 we therefore obtain
@@
Invoking the orthogonality of $f_{\perp}$ and $(\mathbb{P}_{\textup{in}}-\mathbb{P}_{\textup{out}})\mathbf{k}\in\mathcal{L}$, the reproducing property representations 52, and the vector sub-Gaussianity condition 49, we find that
@@ B.1 Subsampling
$\mathbb{E}[\boldsymbol{p}_{\textup{out}}\boldsymbol{p}_{\textup{out}}^{\top}]=\frac{1}{n_{\textup{out}}}\mathop{\mathrm{missing}}{diag}(\boldsymbol{p}_{\textup{in}})+\frac{n_{\textup{in}}(n_{\textup{out}}-1)}{n_{\textup{out}}(n_{\textup{in}}-1)}(\boldsymbol{p}_{\textup{in}}\boldsymbol{p}_{\textup{in}}^{\top}-\frac{1}{n_{\textup{in}}}\mathop{\mathrm{missing}}{diag}(\boldsymbol{p}_{\textup{in}}))=\frac{1}{n_{\textup{out}}}(\frac{n_{\textup{in}}-n_{\textup{out}}}{n_{\textup{in}}-1})\mathop{\mathrm{missing}}{diag}(\boldsymbol{p}_{\textup{in}})+\frac{n_{\textup{in}}(n_{\textup{out}}-1)}{n_{\textup{out}}(n_{\textup{in}}-1)}\boldsymbol{p}_{\textup{in}}\boldsymbol{p}_{\textup{in}}^{\top}.$ (60)
@@
$\mathbb{E}[\operatorname{MMD}^{2}_{\mathbf{K}}(\boldsymbol{p}_{\textup{in}},\boldsymbol{p}_{\textup{out}})]$ $=\boldsymbol{p}_{\textup{in}}^{\top}\mathbf{K}\boldsymbol{p}_{\textup{in}}-2\boldsymbol{p}_{\textup{in}}^{\top}\mathbf{K}\mathbb{E}[\boldsymbol{p}_{\textup{out}}]+\mathbb{E}[\boldsymbol{p}_{\textup{out}}^{\top}\mathbf{K}\boldsymbol{p}_{\textup{out}}]=\mathop{\mathrm{missing}}{tr}(\mathbf{K}\mathbb{E}[\boldsymbol{p}_{\textup{out}}\boldsymbol{p}_{\textup{out}}^{\top}])-\boldsymbol{p}_{\textup{in}}^{\top}\mathbf{K}\boldsymbol{p}_{\textup{in}}$ (61)
$=\frac{1}{n_{\textup{out}}}(\frac{n_{\textup{in}}-n_{\textup{out}}}{n_{\textup{in}}-1})(\mathop{\mathrm{missing}}{tr}(\mathbf{K}\mathop{\mathrm{missing}}{diag}(\boldsymbol{p}_{\textup{in}}))-\boldsymbol{p}_{\textup{in}}^{\top}\mathbf{K}\boldsymbol{p}_{\textup{in}})=\frac{1}{n_{\textup{out}}}(\frac{n_{\textup{in}}-n_{\textup{out}}}{n_{\textup{in}}-1})C_{\mathbf{K}}.$ (62)
@@
and invoke the initial result 62 to conclude.

### v7

@@ 4.2 Thinning attention in practice
To gauge the practical effectiveness of Alg. 1, we recreate the benchmark Tokens-To-Token Vision Transformer (T2T-ViT) and BigGAN image generation experiments of Zandieh et al. (2023). In the T2T-ViT experiment, attention approximations are scored on their ImageNet classification accuracy and computational expense when used as drop-in replacements for the two most expensive attention layers in a pretrained T2T-ViT neural network (Yuan et al., 2021). In the BigGAN experiment, approximations are scored on their computational expense and two popular measures of image generation quality, the Frechet Inception Distance (FID, Heusel et al., 2017) and Inception Score (IS, Salimans et al., 2016). Using the exact implementations and settings provided by Zandieh et al. (2023), we benchmark our PyTorch implementation of Thinformer against exact attention and four leading attention approximations: Performer (Choromanski et al., 2021), Reformer (Kitaev et al., 2020), ScatterBrain (Chen et al., 2021), and KDEformer.
In Tab. 3, we find that Thinformer ($\mathfrak{g}=2$) provides the highest Top-111 accuracy on the ImageNet 2012 validation set (Russakovsky et al., 2015), while running faster than all of the alternatives. In Tab. 4, Thinformer ($\mathfrak{g}=2$) yields better FID and IS than all of the alternatives while running significantly faster than exact, KDEformer, Reformer, and ScatterBrain. Performer runs faster still but at the expense of substantially worse FID and IS. The final attention call of Thinformer can also be combined with optimized attention implementations like FlashAttention (Dao et al., 2022; Dao, 2024) to further reduce the time and memory footprint. We provide PyTorch code replicating this experiment at https://github.com/microsoft/thinformer and supplementary experiment details in Sec. L.1.
@@ 5 Faster SGD Training
We now turn to a second application, accelerating training through gradient reordering. To train a machine learning model parameterized by $\boldsymbol{w}\in\mathbb{R}^{d}$, a standard approach is to minimize the empirical risk $f(\boldsymbol{w})\triangleq\frac{1}{n}\sum_{i=1}^{n}f_{i}(\boldsymbol{w})$ using stochastic gradient descent (SGD) updates,
@@
[…]
@@
Figure 1: Train and test convergence trajectories for mortgage classification with reordered SGD variants. We display mean values ±1$\pm$ 1± 1 standard deviation across 555 random seeds. See Sec. 5.2 for more details.
@@ 5.1 Bridging the dimension gap
Suppose that, for all $i\in[n]$ and $\boldsymbol{w},\boldsymbol{v}\in\mathbb{R}^{d}$, the losses $f$ and $f_{i}$ satisfy
[…]
$f(\boldsymbol{w})-f^{\star}\leq\frac{1}{2\mu}\|{\nabla f(\boldsymbol{w})}\|_{2}^{2}\ \ \textbf{{(PL)}}\text{ for }f^{\star}\triangleq\inf_{\boldsymbol{v}\in\mathbb{R}^{d}}f(\boldsymbol{v}).$ (29)
Then, with probability at least $\frac{1}{2}$, SGD 26 with LKH$(\frac{1}{2K})$ reordering (Alg. 2) and step size $\alpha$ given in App. G satisfies
$\textstyle f($ $\boldsymbol{w}_{K})-f^{\star}\leq\textstyle\widetilde{O}(\frac{r}{n^{2}K^{2}})\quad\text{for}\quad\qquad\qquad$ (30)
$r\triangleq\textstyle\max_{k\in[K]}$ $\ \textstyle\mathrm{rank}_{\epsilon_{k}}([\boldsymbol{x}_{1}^{k},\dots,\boldsymbol{x}_{n}^{k}]),\ \ \bar{\boldsymbol{x}}^{k}\triangleq\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{x}_{i}^{k},$ (31)
$\text{and}\ \ \epsilon_{k}\triangleq\max_{i\in[n]}$ $\ \textstyle\frac{\sqrt{9e\log(4Kn\log(en/2))\log(4Kn)}\|{\boldsymbol{x}_{i}^{k}-\bar{\boldsymbol{x}}^{k}}\|_{2}}{\sqrt{n}}.$ (32)
The proof of Thm. 3 in App. G simply uses Thm. 1 to bound the thinning quality of LKH$(\frac{1}{2K})$ and then adapts the prior SGD analysis of Cooper et al. (2023). Notably, the standard practice of random reshuffling, i.e., SGD with uniform reordering, can only guarantee a significantly slower $\Omega(\frac{1}{nK^{2}})$ rate under these assumptions (Rajput et al., 2020, Thm. 2), while Lu et al. (2022, Thm. 4) implies a similar but dimension-dependent $\widetilde{O}(\frac{d}{n^{2}K^{2}})$ rate for SBW reordering. Thm. 3 matches the minimax lower bound of Cha et al. (2023, Thm. 4.5) up to the $\epsilon$-rank parameter and shows that dimension dependence can be avoided when the gradient update matrices $[\boldsymbol{x}_{1}^{k},\dots,\boldsymbol{x}_{n}^{k}]$ are low-rank, or, more generally, $\epsilon=O(\frac{\log(Kn)}{\sqrt{n}})$-approximable by low-rank matrices.
@@ 5.2 Bridging the theory-practice gap
Two criticisms levied by Lu et al. (2022) against the SBW algorithm were the need to estimate the maximum Euclidean norm of any possible gradient vector in advance and the need to tune its free hyperparameter. LKH$(\frac{1}{2K})$ has neither of these drawbacks as it automatically adapts to the scale of each input and has no hyperparameters to tune. Moreover, with a linear kernel, LKH$(\frac{1}{2K})$ can be run online in $O(nd)$ time. Hence, LKH$(\frac{1}{2K})$ is a promising substitute for the greedy thinning of Lu et al. (2022); Cooper et al. (2023). Indeed, when we recreate the Home Mortgage Disclosure Act logistic regression experiment of Cooper et al. (2023) with a single worker (Fig. 1), we find that LKH-SGD strongly outperforms the standard practice of random reshuffling (RR) and the theoretically justified but overly conservative CD-GraB: SBW variant. In addition, LKH-SGD matches the state-of-the-art test accuracy of CD-GraB: Greedy and lags only slightly in terms of training convergence.
The accelerated convergence rate of LKH-SGD over the standard slow SGD rate of RR provides a direct verification of the Thm. 3 guarantee, and we further verify in Fig. L.1 that the singular values of the gradient update matrices drop off steeply, resulting in relatively small $\epsilon_{k}$-ranks (see Fig. L.1). See https://github.com/microsoft/khsgd for PyTorch code replicating this experiment and Sec. L.2 for supplementary experiment details.
@@ 6 Cheap Two-Sample Testing
Our final application is two-sample testing, determining whether two datasets are drawn from the same underlying distribution. We observe independent samples $\mathcal{X}\triangleq(\boldsymbol{x}_{i})_{i=1}^{m}$ and $\mathcal{Y}\triangleq(\boldsymbol{y}_{j})_{j=1}^{n}$ from the unknown distributions $\mathbb{P}$ and $\mathbb{Q}$ respectively, and we seek to accept or reject the null hypothesis that $\mathbb{P}=\mathbb{Q}$. Standard kernel MMD tests tackle this task by computing the empirical MMD
@@ 6.1 Low-rank analysis of Compress Then Test
$O(4^{\mathfrak{g}}(m+n)(s+\log_{4}\mathopen{}\mathclose{{\left(\frac{m+n}{s}-\mathfrak{g}}}\right)))$ (35)
@@
Instantiate the assumptions of Thm. 4 with $\mathbf{k}=\mathbf{k}_{\textup{deep}}$ 40. If the inputs $(\phi(\boldsymbol{x}_{1}),\boldsymbol{x}_{1},\phi(\boldsymbol{y}_{1}),\boldsymbol{y}_{1})$ are sub-Gaussian, that is,
@@
Cors. 3 and 4 follow from explicitly bounding the eigenvalues of the generated deep kernel matrices as in 17 and 20. By Kim & Schrab (2023, Thm. 8), the separation rate of Thm. 4 is minimax optimal up to the inflation factor $\mathbsf{R}_{\mathbf{k}}/2^{\mathfrak{g}}$ and, hence, those of Cors. 3 and 4 are minimax optimal up to log factors.
One could alternatively bound the compression error of KT-Compress$(\delta)$ using the covering number approach of Dwivedi & Mackey (2022, Thm. 2, Prop. 3). In the setting of Cor. 3, the argument of App. J combined with this distinct analysis would yield an alternative error inflation factor $\mathbsf{\tilde{R}}_{\mathbf{k}_{\textup{deep}}}/2^{\mathfrak{g}}$ with worse dimension dependence,
@@ 6.2 Powerful deep kernel testing in near-linear time
Figure 2: Time-power trade-off curves for detecting Higgs bosons with deep kernel MMD tests. We plot mean values ±1$\pm$ 1± 1 standard error across 100010001000 independent trials with level $\alpha=0.05$ and $\mathcal{B}=100$ permutations.
@@
Fig. 2 compares the time-power trade-off curves induced by three fast kernel testing approaches to this problem: Subsampling, a standard wild-bootstrap MMD test (Chwialkowski et al., 2014) that simply evaluates empirical $\operatorname{MMD}_{\mathbf{k}_{\textup{deep}}}$ using $n_{\textup{out}}=m_{\textup{out}}$ uniformly subsampled points; W-Block, a wild-bootstrap test that averages $\frac{n}{B}$ subsampled squared $\operatorname{MMD}_{\mathbf{k}_{\textup{deep}}}$ estimates based on $n_{\textup{out}}=m_{\textup{out}}=B$ points (Zaremba et al., 2013); and CTT with $s=32$ bins and varying $\mathfrak{g}$. We find that the CTT curve dominates that of the alternative methods and matches the power of an exact MMD test (Subsampling with $n_{\textup{out}}=n$) in a fraction of the time. The improvements of CTT over the standard power-runtime trade-off of Subsampling provides a direct verification of the Thm. 4 guarantee, and we additionally verify in Fig. L.2 that the empirical inflation factor $\widehat{\mathbsf{R}}_{\mathbf{k}_{\textup{deep}}}=O(\log^{5}(n))$ in this setting due to approximate low-rankness. See https://github.com/microsoft/deepctt for PyTorch code replicating this experiment and Sec. L.3 for supplementary experiment details.
@@ Appendix A Appendix Notation and Definitions
If, in addition, the event has probability 111, we say that $\boldsymbol{w}$ is $(\mathbf{K},\nu)$-sub-Gaussian.
Notably, a thinning algorithm is $(\mathbf{K},\nu,\delta)$-sub-Gaussian if and only if its associated vector $\boldsymbol{p}_{\textup{in}}-\boldsymbol{p}_{\textup{out}}$ is $(\mathbf{K},\nu)$-sub-Gaussian on an event $\mathcal{E}$ of probability at least $1-\delta/2$.
@@
If, in addition, the event has probability 111, we say that $\phi$ is $(\mathbf{k},\nu)$-sub-Gaussian.
@@
Invoking the representations 49 and the functional sub-Gaussianity condition 47 we therefore obtain
@@
Invoking the orthogonality of $f_{\perp}$ and $(\mathbb{P}_{\textup{in}}-\mathbb{P}_{\textup{out}})\mathbf{k}\in\mathcal{L}$, the reproducing property representations 49, and the vector sub-Gaussianity condition 46, we find that
@@ B.1 Subsampling
$\mathbb{E}[\boldsymbol{p}_{\textup{out}}\boldsymbol{p}_{\textup{out}}^{\top}]=\frac{1}{n_{\textup{out}}}\mathop{\mathrm{diag}}(\boldsymbol{p}_{\textup{in}})+\frac{n_{\textup{in}}(n_{\textup{out}}-1)}{n_{\textup{out}}(n_{\textup{in}}-1)}(\boldsymbol{p}_{\textup{in}}\boldsymbol{p}_{\textup{in}}^{\top}-\frac{1}{n_{\textup{in}}}\mathop{\mathrm{diag}}(\boldsymbol{p}_{\textup{in}}))=\frac{1}{n_{\textup{out}}}(\frac{n_{\textup{in}}-n_{\textup{out}}}{n_{\textup{in}}-1})\mathop{\mathrm{diag}}(\boldsymbol{p}_{\textup{in}})+\frac{n_{\textup{in}}(n_{\textup{out}}-1)}{n_{\textup{out}}(n_{\textup{in}}-1)}\boldsymbol{p}_{\textup{in}}\boldsymbol{p}_{\textup{in}}^{\top}.$ (57)
@@
$\mathbb{E}[\operatorname{MMD}^{2}_{\mathbf{K}}(\boldsymbol{p}_{\textup{in}},\boldsymbol{p}_{\textup{out}})]$ $=\boldsymbol{p}_{\textup{in}}^{\top}\mathbf{K}\boldsymbol{p}_{\textup{in}}-2\boldsymbol{p}_{\textup{in}}^{\top}\mathbf{K}\mathbb{E}[\boldsymbol{p}_{\textup{out}}]+\mathbb{E}[\boldsymbol{p}_{\textup{out}}^{\top}\mathbf{K}\boldsymbol{p}_{\textup{out}}]=\mathop{\mathrm{tr}}(\mathbf{K}\mathbb{E}[\boldsymbol{p}_{\textup{out}}\boldsymbol{p}_{\textup{out}}^{\top}])-\boldsymbol{p}_{\textup{in}}^{\top}\mathbf{K}\boldsymbol{p}_{\textup{in}}$ (58)
$=\frac{1}{n_{\textup{out}}}(\frac{n_{\textup{in}}-n_{\textup{out}}}{n_{\textup{in}}-1})(\mathop{\mathrm{tr}}(\mathbf{K}\mathop{\mathrm{diag}}(\boldsymbol{p}_{\textup{in}}))-\boldsymbol{p}_{\textup{in}}^{\top}\mathbf{K}\boldsymbol{p}_{\textup{in}})=\frac{1}{n_{\textup{out}}}(\frac{n_{\textup{in}}-n_{\textup{out}}}{n_{\textup{in}}-1})C_{\mathbf{K}}.$ (59)
@@
and invoke the initial result 59 to conclude.